If Harry can do a job in 2 days, and Sally can do the job in 4 days, and Henry can do the job in 6 days how long it will take to do the job if they all work on it at the same time?
How would you set up the equation to solve this problem?

Answer by stanbon(57424)

If Harry can do a job in 2 days, and Sally can do the job in 4 days, and Henry can do the job in 6 days how long it will take to do the job if they all work on it at the same time?
---
Harry rate is 1/2 job/day
Sally rate is 1/4 job/day
Henry rate is 1/6 job/day
---
Together rate is 1/x job/day
-----------------------------------
Equation:
rate + rate + rate = together rate
---
1/2 + 1/4 + 1/6 = 1/x
----
6x + 3x + 2x = 12
11x = 12
x = 12/11 = 1 1/11 days (time reqrd to do the job together)
================================
Cheers,
Stan H.